Angie’s Bake Shop makes birthday chocolate chip cookies that cost $2 each. Angie expects that 10% of the cookies will crack and be discarded. Angie wants a 60% markup on cost and produces 100 cookies. Assume that Angie can sell the cracked cookies for $1.10 each. What should Angie price each cookie? (Round your answer to the nearest cent.)

I kept coming up with $3.56 but it's not correct. please show work
Answer #1

Given cost of single cookie = $2

total cookie produced = $100

ideally if all cookies are sold at a 60% markup on cost, total revenue from selling = 2*100*1.6 = $320

Now, 10% of cookies cracked and sold at $1.1 each.

So, revenue from 10 cookies = 1.1* 10 = $11

remaining sales should be recovered from the 90 cookies, so they must produce a revenue of 320-11 = $309

so, selling price per cookie = 309/90 = $3.43 per cookie